OpenStudy (anonymous):

Show the expanded form and simplification of x^4/x^2. Explain in your own words how you can simplify x^4/x^2 without having to write the expanded form. Part two:Part 2: Create your own fraction with like bases, coefficients, and show its simplification.

OpenStudy (hba):

See this, \[\huge\frac{ a^3 }{ a }=a^{3-1}=a^2\]

OpenStudy (anonymous):

yes ?

OpenStudy (hba):

Now you can solve your own ques, \[\huge\frac{ x^4 }{ x^2 }=..\]

OpenStudy (hba):

@KCB Start working.

OpenStudy (hba):

@singhmmm Giving direct answers is not allowed on OS Please check the CoC- http://openstudy.com/code-of-conduct

OpenStudy (hba):

@singhmmm I think i told you something

OpenStudy (hba):

Please let the asker participate in the solution rather than you doing all the work

OpenStudy (anonymous):

k